# CSF: Fixed-outline Floorplanning Based on the Conjugate Subgradient Algorithm Assisted by Q-Learning

### v1

@@ V-B Performance comparison of legalization algorithms
Figure 1: The Floorplans of GSRC benechmarks generated by CSF-qq.
VI Conclusion
The state-of-the-art researches demonstrate that the analytic floorplanning algorithms generally outperform the counterparts based on the working principle of combinatorial optimization. These analytic floorplanning algorithms typically employ some gradient-based programming algorithms addressing the smoothed objectives, which introduces inevitable distortion to the characteristics of the mathematical model. Consequently, the generated floorplan sometimes is not as optimal as expected.
In this paper, we propose to solve the original nonsmooth objective of floorplanning by the CSA, and the Q-learning strategy is introduced to improve the convergence performance of the CSA-based floorplanning algorithm. Experimental results show that the proposed CSF accelerated by Q-learning (CSF-qq) is competitive to the state-of-the-art fixed-ouline floorplanning algorithms in the case that only hard modules are considered. However, the Q-learning strategy leads to a slight increase on the runtime. Our future work will try to decrease the time complexity of the Q-learning accelerated algorithm, and extend it to the case of large-scale floorplanning with both soft and hard modules.

### v2

@@ 5.3.1 Floorplanning within compact layout regions
Figure 2: The Floorplans of GSRC benechmarks generated by CSF-$q$.
Table 7: Performance Comparison for the Floorplanning Restricted in Layout Regions with 15% Whitespace
Circuit Method HPWL($\mu m$) CPU($s$) SR($\%$) IR($\%$)
ami33 Parquet-4.5 78650 0.16 60 -6.07
FFA-CD - - - -
CSF-$q$ 83427 0.91 83.3 -
ami49 Parquet-4.5 962654 0.37 60 5.93
FFA-CD - - - -
CSF-$q$ 905566 0.58 100 -
n100 Parquet-4.5 334719 1.73 100 13.31
FFA-CD 293578 0.70 100 1.17
CSF-$q$ 290156 1.76 100 -
n200 Parquet-4.5 620097 7.78 100 16.27
FFA-CD 521140 1.95 100 0.37
CSF-$q$ 519211 5.35 100 -
n300 Parquet-4.5 768747 16.86 100 23.39
FFA-CD 588118 3.44 100 -0.14
CSF-$q$ 588951 12.40 100 -
The numerical results show that CSF-qq is generally competitive with the compared algorithms. Compared to Parquet-4.5, the wirelength on GSRC benchmarks is improved by $13.31\sim 23.39\%$, the running time is reduced to a large extent. In comparison with FFA-CD, the CSF-qq generates floorplans on n100 and n200 with shorter wirelength, and the wirelength of problem n300 is a bit worse than FFA-CD by -0.14%. Since a population-based CSA framework assisted by Q-learning is employed by CSF-qq, its running time on GSRC benchmarks is slightly longer than FFA-CD.
Meanwhile, the global convergence of CSF-qq is guaranteed by the optimizer CSAQ. It can address the MCNC benchmarks better, achieving 100% success rate (SR) on ami33 and ami49. In comparison, the SR of Parquet-4.5 is 60%, and there is no reference data available for FFA-CD in Ref.[22].
5.3.2 Wirelength minimization within spacious layout regions
To validate the performance of CSF-qq on minimization of wirelength, we compare it with two representative algorithms, the Fast-SA and the Per-RMAP, where the fixed-outlines of layout regions are generated according to setting $s2$ presented in Tab. 2 [17]. As presented in Tab. 8, CSF-qq outperforms both Fast-SA and Per-RMAP on the GSRC benchmarks in terms of wirelength, however, it gets slightly longer wirelength when addressing the MCNC benchmarks. With regard to the running time, both CSF-qq and Per-RMAP beat Fast-SA markedly. Note that the results of Per-RMAP are cited from Ref. [17] where it is implemented on a server equipped with 2-way Intel Xeon Gold 6248R@3.0-GHz CPUs and 768-GB DDR4-2666-MHz memory, CSF-qq is also competitive to Per-RMAP in terms of running time.
Table 8: Performance Comparison on [17] Fixed-outline Floorplanning
Circuit Method HPWL($\mu m$) CPU($s$) SR($\%$) IR($\%$)
ami33 Fast-SA - - - -
Per-RMAP 63079 0.14 - -2.82
CSF-$q$ 64857 0.14 100 -
ami49 Fast-SA - - - -
Per-RMAP 689296 1.70 - -5.56
CSF-$q$ 727610 0.27 100 -
n100 Fast-SA 287646 10.72 - 3.22
Per-RMAP 282596 1.01 - 1.49
CSF-$q$ 278374 1.74 100 -
n200 Fast-SA 516057 69.86 - 2.29
Per-RMAP 518722 2.93 - 2.79
CSF-$q$ 504258 3.75 100 -
n300 Fast-SA 603811 133.63 - 3.05
Per-RMAP 626061 4.11 - 6.49
CSF-$q$ 585419 5.99 100 -
6 Conclusion
The state-of-the-art researches demonstrate that the analytic floorplanning algorithms generally outperform the counterparts based on the working principle of combinatorial optimization. These analytic floorplanning algorithms typically employ some gradient-based optimization algorithms addressing the smoothed objectives, which introduces inevitable distortion to the characteristics of the mathematical model. Consequently, the generated floorplan sometimes is not as optimal as expected.
In this paper, we propose to address the original nonsmooth optimization model of floorplanning by the CSA, and the Q-learning strategy is introduced to improve the convergence performance of the CSA-based floorplanning algorithm. Experimental results show that the proposed CSF accelerated by Q-learning (CSF-qq) is competitive to the state-of-the-art fixed-outline floorplanning algorithms in the case that only hard modules are considered. However, the population-based iteration scheme leads to a slight increase runtime in some investigated scenarios, and its performance on small-scale MCNC benchmarks is not as good as that on GSRC problems. Accordingly, our future work will focus on the development of orientation optimization of modules, try to decrease the time complexity of the Q-learning accelerated algorithm, and extend it to the case of large-scale floorplanning with both soft and hard modules.